integrate the following integral
∫ x/sqrt(x + 4) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{x}{\sqrt{x+4}}dx
I = \int \frac{x+4-4}{\sqrt{x+4}}dx
I = \int [\sqrt{x+4}-\frac{4}{\sqrt{x+4}}]dx
I = \int \sqrt{x+4}dx-\int \frac{4}{\sqrt{x+4}}dx
I = \frac{2(x+4)^{3/2}}{3}-\int \frac{4}{\sqrt{x+4}}dx
I = \frac{2(x+4)^{3/2}}{3}-8\sqrt{x+4}+c